The question is:
You are given f(x)=-x+6 and g(x)=f(x+3), write an equation for g.
This question was on my homework and I am completely confused. Please help.
Thanks.

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

You are given f(x)=-x+6 and g(x)=f(x+3), write an equation for g.
--------
Not an equation for g, for g(x)
-----
Find f(x+3)
Wherever you see x, sub (x+3)
f(x+3) = -(x+3) + 6 = -x + 3
g(x) = -x + 3

Answer by ikleyn(52800)

.

Wherever you see x, sub (x+3)


f(x+3) = -(x+3) + 6 = -x + 3.


g(x) = -x + 3.        ANSWER